OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (anonymous):

Ting tong? Anybody here?

OpenStudy (anonymous):

Slope for equation is Standard form ie \(ax + by = c\) is given by: \[Slope = \frac{- x \; \; coefficient}{y \; \; coefficient}\]

OpenStudy (anonymous):

m=-4/5, b=-9/5